OpenStudy (barreraa):

Simplify. –15 – (–3) + 4 (Points : 4) –16 –8 –22 8

OpenStudy (barreraa):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

-15+3+4=-12+4=????

OpenStudy (alexandervonhumboldt2):

- - =+

OpenStudy (barreraa):

-8

OpenStudy (alexandervonhumboldt2):

yes